OpenStudy (anonymous):

1/9[8Ln(x+8)-Lnx-Ln(x^2-49)]

OpenStudy (anonymous):

What about that? Do you have to write that in a shorter form? Do you have to evaluate the expression for a particular x?

OpenStudy (mertsj):

\[\frac{1}{9}[8\ln (x+8)-(\ln x+\ln (x^2-49))]=\frac{1}{9}[8\ln (x+8)-\ln x(x^2-49)]\] \[\frac{1}{9}[\ln (x+8)^8-\ln (x^3-49x)]=\frac{1}{9}\ln \frac{(x+8)^8}{x^3-49x}\] \[\ln (\frac{x+8)^8}{x^3-49x})^{\frac{1}{9}}\]